# Data Bias According to Bipol: Men are Naturally Right and It is the Role of Women to Follow Their Lead

### v1

@@ Abstract
Caution: This paper contains examples of what some may consider as stereotypes or offensive text.
@@ 1 Introduction
The problem of social bias in data is a pressing one. Recent news about social bias of artificial intelligence (AI)systems, such as Alexa111bbc.com/news/technology-66508514 and ChatGPT,222bloomberg.com/news/newsletters/2022-12-08/chatgpt-open-ai-s-chatbot-is-spitting-out-biased-sexist-results shows that the age-old problem persists with data, which is used to train machine learning (ML)models. Social bias is the inclination or prejudice for, or against, a person, group or idea, especially in a way that is considered to be unfair, which may be based on race, religion or other factors (Bellamy et al., 2018; Antoniak & Mimno, 2021; Mehrabi et al., 2021; Alkhaled et al., 2023). It can also involve stereotypes that generalize behavior to groups (Brownstein, 2019). It can unfairly skew the output of ML models (Klare et al., 2012; Raji et al., 2020). Languages with fewer resources than English are also affected (Rescigno et al., 2020; Chávez Mulsa & Spanakis, 2020; Kurpicz-Briki, 2020). For example, in Italian, the female gender is under-represented due to the phenomena such as the “inclusive masculine" (when the masculine is over-extended to denote groups of both male and female referents) (Luccioli et al., ; Vanmassenhove & Monti, 2021).
In this work, we are motivated to address the research question of how much bias exists in the text data of multiple languages, if at all bias exists in them. We particularly investigate 6 benchmark datasets on the English GLUE/SuperGLUE leaderboards (Wang et al., 2018; 2019) and one dataset each for the other 4 languages: Italian, Dutch, German, and Swedish. First, we train SotA multilingual Text-to-Text Transfer Transformer (mT5)(Xue et al., 2021) and multilingual Bidirectional Encoder Representations from Transformers (mBERT) models for bias classification on the multi-axes bias dataset (MAB)(Alkhaled et al., 2023) for each language. For the evaluations, we search through the literature to compare different metrics or evaluation methods as shown in Table 1 and discussed in Section 2. This motivates our choice of bipol, the multi-axes bias metric, which we then compare in experiments with a lexica baseline method. In addition, to confirm the unverified assumption that toxic comments contain bias (Sap et al., 2020; Alkhaled et al., 2023), we annotate 200 randomly-selected samples from the training set of the English MAB.
@@ 2 Literature Review
Although English usually gets more support and attention in the literature, there have been attempts at measuring and mitigating bias in other languages. Testing for the presence of bias in Italian often has a contrastive perspective with English, with a focus on gender bias (Gaido et al., 2021; Rescigno et al., 2020). MuST-SHE (Bentivogli et al., 2020) and gENder-IT (Vanmassenhove & Monti, 2021) are examples of gender bias evaluation sets. Going beyond gender bias, Kurpicz-Briki & Leoni (2021) and Huang et al. (2020) also identified biases related to people’s origin and speakers’ age. It is essential to remember that the mentioned biases can be vehicles for misogynous and hateful discourse (El Abassi & Nisioi, 2020; Attanasio et al., 2022; Merenda et al., 2018).
@@
Bias studies for Dutch mostly consider binary gender bias. Chávez Mulsa & Spanakis (2020) investigate gender bias in Dutch static and contextualized word embeddings by creating Dutch versions of the Word/Sentence Embedding Association Test (WEAT/SEAT) (Caliskan et al., 2017; May et al., 2019). WEAT measures bias in word embeddings and can be limited in scope, in addition to having sensitivity to seed words. McCurdy & Serbetci (2020) perform a similar evaluation in a multilingual setup to compare the effect of grammatical gender saliency across languages. Several works use different NLP techniques to evaluate bias in corpora of Dutch news articles (Wevers, 2019; Kroon et al., 2020; Kroon & van der Meer, 2021; Fokkens et al., 2018) and literary texts (Koolen & van Cranenburgh, 2017).
@@
In Kurpicz-Briki (2020), bias is measured with regards to place of origin and gender in German word embeddings using WEAT (Caliskan et al., 2017). In Kurpicz-Briki & Leoni (2021), an automatic bias detection method (BiasWords) is presented, through which new biased word sets can be identified by exploring the vector space around the well-known word sets that show bias. In the template-based study of Cho et al. (2021), on gender bias in translations, the accuracy of gender inference was measured for multiple languages including German. It was found that, particularly for German, the inference accuracy and disparate impact were lower for female than male, implying that certain translations were wrongly performed for cases that required female inference. Since German is a grammatically gendered, morphologically rich language, Gonen & Goldberg (2019) found that debiasing methods of Bolukbasi et al. (2016) were ineffective on german word embeddings.
@@
For Swedish, the main focus of bias research appears to be on gender. Sahlgren & Olsson (2019) show with their experiments that gender bias is present in pretrained Swedish language models. Katsarou et al. (2022) and Precenth (2019) found that the male gender tends to be associated with higher-status professions. A study with data from mainstream news corpora by Devinney et al. (2020) shows that women are associated with concepts like family, communication and relationships.
@@ 3 Bipol
The bipol metric uses a two-step mechanism for estimating bias in text data: binary classification and sensitive term evaluation using lexica. It has maximum and minimum values of 1 and 0, respectively. Bipol is expressed in Equations 1b and 1c from the main Equation 1a, where $b_{c}$ is the classification component and $b_{s}$ is the sensitive term evaluation component. Details about the metric are discussed in Alkhaled et al. (2023).
@@
In step 1, a trained model is used to classify all the samples. The ratio of the biased samples to the total samples predicted is determined. The tp, fp, tn, and fn are values of the true positives, false positives, true negatives, and false negatives, respectively. Since there’s hardly a perfect classifier, the positive error rate is usually reported. False positives are known to exist in similar classification systems like spam detection and automatic hate speech detection (Heron, 2009; Feng et al., 2018; Adewumi et al., 2023b).
Step 2 is similar to term frequency-inverse document frequency (TF-IDF) in that it is based on term frequency (Salton & Buckley, 1988; Ramos et al., 2003), Biased samples from step 1 are evaluated token-wise along all possible bias axes, using all the lexica of sensitive terms. An axis is a domain such as gender or race. Tables 2 and 3 provide the lexica sizes. For English and Swedish, we use the same lexica released by Alkhaled et al. (2023) and Adewumi et al. (2023c), respectively. For the other 3 languages, we create new lexica of terms (e.g. she & her) associated with specific gender or stereotypes from public sources.444fluentu.com/blog/italian/italian-nouns, en.wiktionary.org/wiki/Category:Italian_offensive_terms, Dutch_profanity, Category:German_ethnic_slurs Some of the terms in the lexica were selected from the sources based on the topmost available. These may also be expanded as needed, since bias terms are known to evolve (Haemmerlie & Montgomery, 1991; Antoniak & Mimno, 2021). The non-English lexica are small because fewer terms are usually available in other languages compared to the high-resource English language and we use the same size across the languages to be able to compare performance somewhat.
@@ 4.1.1 MAB
A fitness ranking that drops each month as more and more fat people move to Colorado.
biased
6231681
167
@@
Women ! ! ! PERSON taught you better than that. Shame on you!
biased
6233025
628
@@
"The article itself touches on the topic of abortion, so absolutely it’s appropriate to discuss it."
unbiased
451965
1486981
"Het artikel zelf raakt het onderwerp abortus, dus het is absoluut gepast om het te bespreken."
"Der Artikel selbst berührt das Thema Abtreibung, also ist es absolut angebracht, es zu diskutieren."
"L’articolo stesso tocca il tema dell’aborto, quindi è assolutamente opportuno discuterne."
"Artikeln själv berör ämnet abort, så absolut är det lämpligt att diskutera det."
@@
The Italian, Dutch and German datasets were machine-translated from MAB  [^5] The reference provides details of the annotation of the base data. (Alkhaled et al., 2023) with the Helsinki-NLP model (Tiedemann & Thottingal, 2020). Each translation took about 48 hours on one GPU. Examples from the data are provided in Table 4. Table 5 provides statistics about the datasets. For quality control (QC), we verified translation by back-translating some random samples using Google NMT. Personal identifiable information (PII) was removed from the MAB dataset using the spaCy library. The 3 datasets are used to train new bias classifiers. We also train on the original English and the Swedish.
@@
Culture-specific biases may not be represented in the MAB versions for the translated languages because the original dataset is in English. This is a limitation. However, bias is also a universal concern, such that there are examples that span across cultures. For instance, examples 1 and 2 in Table 4 are of universal concern because individuals with non-conforming bodies and women should be respected, regardless of culture or nationality. Hence, the MAB versions are relevant for bias detection, though they were translated.
@@ 4.1.2 Evaluation datasets
We selected datasets that are available on the HuggingFace (Wolf et al., 2020) Datasets. We evaluated the first 1,000 samples of each training split due to resource constraints. For English, we evaluated the sentence column of Corpus of Linguistic Acceptability (CoLA)(Warstadt et al., 2019), the sentence column of Question-Answering Natural Language Inference (QNLI)(Wang et al., 2018), the sentennce1 column of Microsoft Research Paraphrase Corpus (MRPC)(Dolan & Brockett, 2005), the premise column of Multi-Genre Natural Language Inference (MNLI)(Williams et al., 2018), the premise column of the CommitmentBank (CB)dataset (De Marneffe et al., 2019), and the passage column of Reading Comprehension with Commonsense Reasoning Dataset (ReCoRD)(Zhang et al., 2018). For Italian, we evaluated the context column of the Stanford Question Answering Dataset (SQuAD)(Croce et al., 2018; Rajpurkar et al., 2016); for Dutch, the sentence1 column of the Semantic Textual Similarity Benchmark (STSB)(Cer et al., 2017); for German, the text column of the German News Articles Datasets 10k (GNAD10)(Schabus et al., 2017); for Swedish, the premise of the CB.
@@ 5.1 Error analysis & qualitative results
nicky approached her with the assumption that men are naturally right and it is the role of women to follow their lead. constance whose confidence was growing daily was not prepared to give in to nickys wishes merely because of his sex. if she felt he was right then she agreed with him.
biased
85
b: thats true. a: so. b: uh the other argument is that the death penalty is a deterrent and i really dont uh agree with that. i dont think anyone who would commit uh a crime that would get them the death penalty would stop at the moment and say well i was about to kill and dismember this person but oh if they catch me theyre going to kill me so i better not do it. i just dont think uh that it works that way. a: yeah. i dont think its done.
biased
133
@@
b: yeah and the ground will filter some of it but not all of it. a: no not when you figure i didnt realize one cow produces that much manure
biased
137
@@
if you give him enough opportunity every senator will succumb to corruption.
biased
266
every senator becomes more corrupt the more lobbyists he talks to.
biased
277
@@
bill squeezed the puppet through the hole.
biased
82
@@ 5.2 Consistent prediction with perturbation
An interesting property of relative consistency that we observed with the model predictions, as demonstrated with the CoLA dataset, is that when sentences are perturbed, the model mostly maintains its predictions, as long as the grounds for prediction (in this case - over-generalization) remain the same. The perturbations are inherent in the CoLA dataset itself, as the dataset is designed that way. Some examples are provided in Table 9, where 6 out of 8 are correctly predicted. This property is repeated consistently in other examples not shown here.
comment_text prediction csv id
if you give him enough opportunity every senator will succumb to corruption.
biased
266
you give him enough opportunity and every senator will succumb to corruption.
biased
267
we gave him enough opportunity and sure enough every senator succumbed to corruption.
unbiased
268
if you give any senator enough opportunity he will succumb to corruption.
biased
269
you give any senator enough opportunity and he will succumb to corruption.
biased
270
you give every senator enough opportunity and he will succumb to corruption.
biased
271
we gave any senator enough opportunity and sure enough he succumbed to corruption.
biased
272
we gave every senator enough opportunity and sure enough he succumbed to corruption.
unbiased
273
@@ 5.3 Explainability by graphs
We show explainability by visualization using graphs. Bipol produces a dictionary of lists for every evaluation and we represent 2 in top-5 frequent terms bar graph. Many of the datasets display overall male bias. As examples, figures 4 and 5 show that the GNAD10 and STSB datasets have overall male and female bias, respectively.
@@
Figure 5: Top 5 frequent terms in the STSB dataset (paired terms are only for comparison).
@@ 6 Conclusion
The findings of this work show that bias besets Natural Language Processing (NLP)datasets regardless of language, including benchmark datasets on the GLUE/SuperGLUE leaderboards. We introduced MAB datasets in 3 languages for training models in bias detection. Each has about 2 million labeled samples. We also contribute lexica of bias terms for the languages. In addition, we verified the assumption that toxic comments contain bias. It may be impossible to completely remove bias from data or models, since they reflect the real world, but resources for estimating bias can provide insight into mitigation strategies for reducing bias. Future work may explore ways of minimising false positives in classifiers to make them more effective. One may also explore how this work scales to other languages or how multilignual models compare to language-specific monolingual models or large language models (LLM s). Regarding culture-specific biases in datasets, one solution will be to collect data from the specific cultures/nationalities to capture these biases.
@@ Acknowledgments
This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP), funded by the Knut and Alice Wallenberg Foundation and counterpart funding from Luleå University of Technology (LTU). The authors also thank Björn Backe for his insightful comments during the writing of this paper.

### v2

@@ 1 Introduction
The problem of social bias in data is a pressing one. Recent news about social bias of artificial intelligence (AI)systems, such as Alexa111bbc.com/news/technology-66508514 and ChatGPT,222bloomberg.com/news/newsletters/2022-12-08/chatgpt-open-ai-s-chatbot-is-spitting-out-biased-sexist-results shows that the age-old problem persists with data, which is used to train machine learning (ML)models. Social bias is the inclination or prejudice for, or against, a person, group or idea, especially in a way that is considered to be unfair, which may be based on race, religion or other factors (Bellamy et al., 2018; Antoniak and Mimno, 2021; Mehrabi et al., 2021; Alkhaled et al., 2023). It can also involve stereotypes that generalize behavior to groups (Brownstein, 2019). It can unfairly skew the output of ML models (Klare et al., 2012; Raji et al., 2020). Languages with fewer resources than English are also affected (Rescigno et al., 2020; Chávez Mulsa and Spanakis, 2020; Kurpicz-Briki, 2020). For example, in Italian, the female gender is under-represented due to the phenomena such as the “inclusive masculine" (when the masculine is over-extended to denote groups of both male and female referents) (Luccioli et al., ; Vanmassenhove and Monti, 2021).
In this work, we are motivated to address the research question of how much bias exists in the text data of multiple languages, if at all bias exists in them? We particularly investigate 6 benchmark datasets on the English GLUE/SuperGLUE leaderboards (Wang et al., 2018, 2019) and one dataset each for the other 4 languages: Italian, Dutch, German, and Swedish. First, we train SotA multilingual Text-to-Text Transfer Transformer (mT5)(Xue et al., 2021) and multilingual Bidirectional Encoder Representations from Transformers (mBERT) models for bias classification on the multi-axes bias dataset (MAB)for each language, in a similar setup as Alkhaled et al. (2023). For the evaluations, we search through the literature to compare different metrics or evaluation methods as shown in Table 1 and discussed in Section 2. This motivates our choice of bipol, the multi-axes bias metric, which we then compare in experiments with a lexica baseline method. In addition, to confirm the unverified assumption that toxic comments contain bias (Sap et al., 2020; Alkhaled et al., 2023), we annotate 200 randomly-selected samples from the training set of the English MAB.
@@ 2 Literature Review
Although English usually gets more support and attention in the literature, there have been attempts at measuring and mitigating bias in other languages. Testing for the presence of bias in Italian often has a contrastive perspective with English, with a focus on gender bias (Gaido et al., 2021; Rescigno et al., 2020). MuST-SHE (Bentivogli et al., 2020) and gENder-IT (Vanmassenhove and Monti, 2021) are examples of gender bias evaluation sets. Going beyond gender bias, Kurpicz-Briki and Leoni (2021) and Huang et al. (2020) also identified biases related to people’s origin and speakers’ age. It is essential to remember that the mentioned biases can be vehicles for misogynous and hateful discourse (El Abassi and Nisioi, 2020; Attanasio et al., 2022; Merenda et al., 2018).
@@
Bias studies for Dutch mostly consider binary gender bias. Chávez Mulsa and Spanakis (2020) investigate gender bias in Dutch static and contextualized word embeddings by creating Dutch versions of the Word/Sentence Embedding Association Test (WEAT/SEAT) (Caliskan et al., 2017; May et al., 2019). WEAT measures bias in word embeddings and can be limited in scope, in addition to having sensitivity to seed words. McCurdy and Serbetci (2020) perform a similar evaluation in a multilingual setup to compare the effect of grammatical gender saliency across languages. Several works use different NLP techniques to evaluate bias in corpora of Dutch news articles (Wevers, 2019; Kroon et al., 2020; Kroon and van der Meer, 2021; Fokkens et al., 2018) and literary texts (Koolen and van Cranenburgh, 2017).
@@
In Kurpicz-Briki (2020), bias is measured with regards to place of origin and gender in German word embeddings using WEAT (Caliskan et al., 2017). In Kurpicz-Briki and Leoni (2021), an automatic bias detection method (BiasWords) is presented, through which new biased word sets can be identified by exploring the vector space around the well-known word sets that show bias. In the template-based study of Cho et al. (2021), on gender bias in translations, the accuracy of gender inference was measured for multiple languages including German. It was found that, particularly for German, the inference accuracy and disparate impact were lower for female than male, implying that certain translations were wrongly performed for cases that required female inference. Since German is a grammatically gendered, morphologically rich language, Gonen and Goldberg (2019) found that debiasing methods of Bolukbasi et al. (2016) were ineffective on german word embeddings.
@@
For Swedish, the main focus of bias research appears to be on gender. Sahlgren and Olsson (2019) show with their experiments that gender bias is present in pretrained Swedish language models. Katsarou et al. (2022) and Precenth (2019) found that the male gender tends to be associated with higher-status professions. A study with data from mainstream news corpora by Devinney et al. (2020) shows that women are associated with concepts like family, communication and relationships.
@@ 3 Bipol
For the purpose of this work, we summarize bipol here but details are discussed in Alkhaled et al. (2023). The bipol metric uses a two-step mechanism for estimating bias in text data: binary classification and sensitive term evaluation using lexica. It has maximum and minimum values of 1 and 0, respectively. Bipol is expressed in Equations 1b and 1c from the main Equation 1a, where $b_{c}$ is the classification component and $b_{s}$ is the sensitive term evaluation component.
@@
In step 1, a trained model is used to classify all the samples. The ratio of the biased samples to the total samples predicted is determined. The tp, fp, tn, and fn are values of the true positives, false positives, true negatives, and false negatives, respectively. Since there’s hardly a perfect classifier, the positive error rate is usually reported. False positives are known to exist in similar classification systems like spam detection and automatic hate speech detection (Heron, 2009; Feng et al., 2018).
Step 2 is similar to term frequency-inverse document frequency (TF-IDF) in that it is based on term frequency (Salton and Buckley, 1988; Ramos et al., 2003), Biased samples from step 1 are evaluated token-wise along all possible bias axes, using all the lexica of sensitive terms. An axis is a domain such as gender or race. Tables 2 and 3 provide the lexica sizes. For English and Swedish, we use the same lexica released by Alkhaled et al. (2023) and Adewumi et al. (2023b), respectively. For the other 3 languages, we create new lexica of terms (e.g. she & her) associated with specific gender or stereotypes from public sources.444fluentu.com/blog/italian/italian-nouns, en.wiktionary.org/wiki/Category:Italian_offensive_terms, Dutch_profanity, Category:German_ethnic_slurs Some of the terms in the lexica were selected from the sources based on the topmost available. These may also be expanded as needed, since bias terms are known to evolve (Haemmerlie and Montgomery, 1991; Antoniak and Mimno, 2021). The non-English lexica are small because fewer terms are usually available in other languages compared to the high-resource English language and we use the same size across the languages to be able to compare performance somewhat. The Appendix lists these terms.
@@ 4.1.1 MAB
A fitness ranking that drops each month as more and more fat people move to Colorado. biased 6231681 167
@@
Women ! ! ! PERSON taught you better than that. Shame on you! biased 6233025 628
@@
The Italian, Dutch and German datasets were machine-translated from MAB  [^5] The reference provides details of the annotation of the base data. with the high-quality Helsinki-NLP model (Tiedemann and Thottingal, 2020). Each translation took about 48 hours on one GPU. Examples from the data are provided in Table 4. Table 5 provides statistics about the datasets. For quality control (QC), we verified translation by back-translating some random samples using Google NMT. Personal identifiable information (PII) was removed from the MAB dataset using the spaCy library. The 3 datasets are used to train new bias classifiers. We also train on the original English and the Swedish.
@@
Culture-specific biases may not be represented in the MAB versions for the translated languages because the original dataset is in English. This is a limitation. However, bias is also a universal concern, such that there are examples that span across cultures. For instance, the examples in Table 4 are of universal concern because individuals with non-conforming bodies and women should be respected, regardless of culture or nationality. Hence, the MAB versions are relevant for bias detection, though they were translated.
@@ 4.1.2 Evaluation datasets
We selected datasets that are available on the HuggingFace (Wolf et al., 2020) Datasets. We evaluated the first 1,000 samples of each training split due to resource constraints. The understanding is that if bias is detected in these samples, then scaling over the entire dataset means there’s probability of more bias. For English, we evaluated the sentence column of Corpus of Linguistic Acceptability (CoLA)(Warstadt et al., 2019), the sentence column of Question-Answering Natural Language Inference (QNLI)(Wang et al., 2018), the sentennce1 column of Microsoft Research Paraphrase Corpus (MRPC)(Dolan and Brockett, 2005), the premise column of Multi-Genre Natural Language Inference (MNLI)(Williams et al., 2018), the premise column of the CommitmentBank (CB)dataset (De Marneffe et al., 2019), and the passage column of Reading Comprehension with Commonsense Reasoning Dataset (ReCoRD)(Zhang et al., 2018). For Italian, we evaluated the context column of the Stanford Question Answering Dataset (SQuAD)(Croce et al., 2018; Rajpurkar et al., 2016); for Dutch, the sentence1 column of the Semantic Textual Similarity Benchmark (STSB)(Cer et al., 2017); for German, the text column of the German News Articles Datasets 10k (GNAD10)(Schabus et al., 2017); for Swedish, the premise of the CB.
@@ 5.1 Error analysis & qualitative results
nicky approached her with the assumption that men are naturally right and it is the role of women to follow their lead. constance whose confidence was growing daily was not prepared to give in to nickys wishes merely because of his sex. if she felt he was right then she agreed with him. biased 85
b: thats true. a: so. b: uh the other argument is that the death penalty is a deterrent and i really dont uh agree with that. i dont think anyone who would commit uh a crime that would get them the death penalty would stop at the moment and say well i was about to kill and dismember this person but oh if they catch me theyre going to kill me so i better not do it. i just dont think uh that it works that way. a: yeah. i dont think its done. biased 133
@@
b: yeah and the ground will filter some of it but not all of it. a: no not when you figure i didnt realize one cow produces that much manure biased 137
@@
if you give him enough opportunity every senator will succumb to corruption. biased 266
every senator becomes more corrupt the more lobbyists he talks to. biased 277
@@
bill squeezed the puppet through the hole. biased 82
@@ 5.2 Consistent prediction with perturbation
An interesting property of relative consistency that we observed with the model predictions, as demonstrated with the CoLA dataset, is that when sentences are perturbed, the model mostly maintains its predictions, as long as the grounds for prediction (in this case - over-generalization) remain the same. The perturbations are inherent in the CoLA dataset itself, as the dataset is designed that way. Some examples are provided in Table 9 in the Appendix, where 6 out of 8 are correctly predicted. This property is repeated consistently in other examples not shown here.
@@ 5.3 Explainability by graphs
We show explainability by visualization using graphs. Bipol produces a dictionary of lists for every evaluation and we show the top-5 frequent terms bar graph for the GNAD10 dataset in Figure 4, which has overall male bias. Many of the 10 evaluated datasets display overall male bias.
@@ 6 Conclusion
The findings of this work show that bias besets Natural Language Processing (NLP)datasets regardless of language, including benchmark datasets on the GLUE/SuperGLUE leaderboards. We introduced MAB datasets in 3 languages for training models in bias detection. Each has about 2 million labeled samples. We also contribute lexica of bias terms for the languages. In addition, we verified the assumption that toxic comments contain bias. It may be impossible to completely remove bias from data or models, since they reflect the real world, but resources for estimating bias can provide insight into mitigation strategies for reducing bias. Future work may explore ways of minimizing false positives in classifiers to make them more effective. One may also explore how this work scales to other languages or how multilignual models compare to language-specific monolingual models or large language models (LLM s). Regarding culture-specific biases in datasets, one solution will be to collect data from the specific cultures/nationalities to capture these biases.
@@ Acknowledgments
The authors wish to thank the anonymous reviewers for their valuable feedback. This work is partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP), funded by the Knut and Alice Wallenberg Foundation and counterpart funding from Luleå University of Technology (LTU). The authors also thank Björn Backe for his insightful comments during the writing of this paper.
Appendix A Appendix
Swedish lexica and the English equivalent
  1. 1.
    Gender-female
    i) hon (she) ii) hennes (her) iii) flicka (girl) iv) mor (mother) v) kvinna (woman) vi) dotter (daughter) vii) mormor (grandmother) viii) dam (lady) ix) sondotter (son’s daughter) x) dotterdotter (daughter’s daughter) xi) tjej (girl) xii) tjejer (girls) xiii) gumma (old woman) xiv) fru (wife) xv) slampa (slut) xvi) slyna (slut) xvii) lebb (lesbian)
  2. 2.
    Gender-male
    i) han (he) ii) pojke (boy) iii) kille (boy) iv) far (father) v) farfar (grandfather) vi) hans (his) vii) man (man) viii) son (son) ix) make (husband) x) herre (lord) xi) herrar (lords) xii) sonson (son’s son) xiii) dotterson (daughter’s son) xiv) gubbe (old man) xv) farbro (uncle) xvi) broson (nephew) xvii) män (men)
  3. 3.
    Racial-black
    i) neger (nigger) ii) blåneger (nigger) iii) blåland (nigger) iv) blatte (dark immigrants) v) svartskalle (black head) vi) turk (non-Swedish) vii) sosse (political slur) viii) svarting (blackness) ix) partysvenska (political slur) x) nigga (nigger)
  4. 4.
    Racial-white
    i) svenne (Swedish stereotype) ii) turk (non-Swedish) iii) jugge (derogatory term) iv) sosse (political slur) v) hurrare (Finnish Swedish) vi) lapp (Sami people) vii) Ang mo (red hair) viii) partysvenska (political slur) ix) Ann (White woman) x) rutabaga (Swede)
Example of predictions for the CoLA dataset
comment_text prediction csv id
if you give him enough opportunity every senator will succumb to corruption. biased 266
you give him enough opportunity and every senator will succumb to corruption. biased 267
we gave him enough opportunity and sure enough every senator succumbed to corruption. unbiased 268
if you give any senator enough opportunity he will succumb to corruption. biased 269
you give any senator enough opportunity and he will succumb to corruption. biased 270
you give every senator enough opportunity and he will succumb to corruption. biased 271
we gave any senator enough opportunity and sure enough he succumbed to corruption. biased 272
we gave every senator enough opportunity and sure enough he succumbed to corruption. unbiased 273